Opened 11 years ago

Closed 9 years ago

Last modified 3 years ago

#240 closed flight processing (fixed)

GB09/07, flight day 132/2009, West coast

Reported by: mggr Owned by: knpa
Priority: alpha 4 high Milestone: 2009 Data processing completion
Component: Archiving Keywords:
Cc: knpa, emca, mark1, sbg, peland Other processors: knpa,benj,mark1

Description (last modified by knpa)

Data location: ~arsf/arsf_data/2009/flight_data/uk/GB09_07-2009_132_West_coast

Data arrived from ARSF via SATA disk on 19/May/2009.

Scientific details: plant distribution on dune fields.

Priority: α4L

PI: Andreas Baas

Sensors:

  • Eagle (07/12/09)
  • Hawk (07/12/09)
  • Leica LIDAR (28/10/2009)
  • Digital Photography (15/01/2010)

Change History (59)

comment:1 Changed 11 years ago by mggr

Logsheet missing, runscripts not generated. Requested vectors and logsheet from Gary.

comment:2 Changed 11 years ago by mggr

  • Priority changed from alpha 5 to alpha 4 low

Corrected grading following fixing trac's priority list..

comment:3 Changed 11 years ago by mggr

Note hawk files may need renaming, not in SWIR* form.

comment:4 Changed 11 years ago by mggr

  • Cc mark1 added

comment:5 Changed 11 years ago by mggr

  • Cc sbg peland added

Other points of note from the application:

  • potentially interested in atmospheric correction / level 2 products (adding peland & sbg to cc)
  • no nextmap access, no digimap (Gary will check?)

comment:6 Changed 11 years ago by mggr

Logsheet received, but as a scan in a PDF file. Someone will have to create the run scripts manually.

comment:7 Changed 11 years ago by mggr

Email from PI:

Thanks for the update on the data processing status. I would just like to confirm that we would certainly be interested in atmospheric correction / level 2 products. Also, I have given Gary the OS / Nextmap / Digimap (whatever it's called these days) subscription number of our university, so that should hopefully not be any issue anymore.

Will ask peland to follow up on the atmospheric/level 2 question.

comment:8 Changed 10 years ago by jatt

  • Owner jatt deleted

Project directory copied to ~arsf/workspace.

Applanix has been processed and copied to this location.

All run scripts have been created.

More than one site - project will need to be split into separate project directories.

comment:9 Changed 10 years ago by knpa

  • Owner set to knpa
  • Status changed from new to assigned

comment:10 Changed 10 years ago by knpa

hawk raw files renamed as per SWIRsh convention

comment:11 Changed 10 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

runscripts edited to SWIRsh convention

comment:12 Changed 10 years ago by knpa

  • Resolution fixed deleted
  • Status changed from closed to reopened

comment:13 Changed 10 years ago by knpa

Directory structure divided into 7 trees according to logsheet groupings: 132a through 132g.

comment:14 Changed 10 years ago by knpa

Incorrect GPS end time on H1 header file (given as 00:00:00). Corrected and updated header file.

comment:15 Changed 10 years ago by knpa

No sync info for h6.

comment:16 Changed 10 years ago by knpa

SCT offsets have been found for 45/52 flightlines. The remaining 7 do not have sufficient wobbles or overlapping fls with wobbles. Will need to wait for vectors (or lidar) before these can be corrected.

comment:17 Changed 10 years ago by mggr

Contacted by PI requesting an update and an estimate of time to delivery. Explained the SCT offset issue and the overall LIDAR processing issues. Offered a potential predelivery of the 45 corrected lines with the proviso of no OS error estimate.

comment:18 Changed 10 years ago by knpa

Have put the majority of the flightlines (excluding lines 5-8) (not yet verified by vectors) lev1 and 3 on pml ftp1 as well as data quality report, readme and screenshot of sites overview.

comment:19 Changed 10 years ago by benj

  • Priority changed from alpha 4 low to alpha 4 high

Have agreed to process this as higher priority because funding for the PhD student intended to look at this (Lily Zhang) runs out soon(ish).

comment:20 Changed 10 years ago by emca

  • Owner changed from knpa to emca
  • Status changed from reopened to new

Starting LIDAR processing

comment:21 Changed 10 years ago by mark1

Vectors have arrived.

comment:22 Changed 10 years ago by emca

  • Owner changed from emca to mark1

LIDAR processing has been finished. Vector data has been received for two sites only (f & g). Emailed PI to offer delivery without checking against vectors.

Delivery directory has been created at
~airborne/workspace/GB09_07-2009_132_West_coast/delivery

Passing ticket to Mark1 for delivery checking.

comment:23 Changed 10 years ago by mark1

Checked ascii files load ok in Grass and look good. Made a few minor edits to read_me. Ready to deliver.

comment:24 Changed 10 years ago by mark1

  • Description modified (diff)

LIDAR data delivered on 28th October 2009

comment:25 Changed 10 years ago by mark1

Processed LIDAR data needs to be rsynced back to main arsf repository, probably best to split the project on the main arsf repository to match the processed project directory structure?

comment:26 Changed 10 years ago by mark1

Have processed the raw RCD camera images and tagged the tiff files.

comment:27 Changed 10 years ago by knpa

  • Other processors set to knpa

comment:28 Changed 10 years ago by knpa

Delivery preparation complete, ready for checking

comment:29 Changed 10 years ago by benj

  • Other processors changed from knpa to knpa,benj

comment:30 Changed 10 years ago by benj

Have delivery checked flight a. Issues so far:

  • Replace data quality report with most up-to-date version
  • Band 232 blank in all hawk files (need to look into this but can deliver this dataset anyway - suspect it's a bandset issue, but band 232 is poor quality anyway, plus should be very similar to band 231)
  • Should mention band 232 issue and Hawk dropping frames (frame grabber card issue) under "Quality of data" in the readme.

comment:31 Changed 10 years ago by benj

Also, level 1 files for delivery should include the flight letter (eg. e132a01.bil, not just e13201.bil)

comment:32 Changed 10 years ago by benj

Actually, don't worry about that for this delivery. Generally it's true, but all the line numbers are different for the various bits of this project so there's not a lot of point in spending the time to do it.

comment:33 Changed 10 years ago by benj

Checking the azgcorr command for flight g hawk line 22 gives the following error in azgcorr:

A UK National Grid projection has been chosen for data that is
 outside the coverage of that projection:  48 N to 62 N and 10 W to 4 E
 The approximate centre of this data set is:  1934672355233165591537245121105912802295402703577532870545838032454086581895847108349688088720759117458960729463204240031593719663501917896233906591097033570989736332795780424822577168766990391955691527327118824677766524783913284925690272373793852668180679655723892736.00 S  2090841113072791560219652599314024076327586797782366537025425377203007704621009079605789662969587471652344465029184481219704709311303332865575410371840520542199889353972652820448798628886650676506023337676581152228735021609379087974166344382137771452130393675695114079533135878132719919721720908435540672512.00 W

The location is, obviously, rubbish, so something somewhere's gone funny with the file. Running hawk line 23 is fine.

comment:34 Changed 10 years ago by benj

Made a few minor readme changes, everything else looks fine. Once the comments above about flight a are fixed and hawk line 22 is investigated this can be rsynced back and delivered.

comment:35 Changed 10 years ago by knpa

SCTs

e h
1 0.07 0.09
2 0.04 0.02
3 0.08 0.08
4 0.07 0.04
5 -0.02 0.04
6 -0.01 0.7
7 0.07 0.07
8 0.02 0.05
9 0.02 0.00
10 0.02 0.07
11 0.03 0.01
12 0.06 0.07
13 -0.01 0.07
14 -0.01 0.06
15 0.07 0.06
16 0.06 0.08
17 0.04 0.06
18 0.06 0.03
19 0.09 0.09
20 0.08 0.05
21 0.07 0.06
22 0.05 0.07
23 0.02 0.09
24 0.07 0.03
25 0.09 0.06
26 0.08 0.04

comment:36 Changed 10 years ago by knpa

  • Owner changed from mark1 to knpa
  • Status changed from new to accepted

Hyperspectral delivered 7th December 09.

comment:37 Changed 10 years ago by knpa

  • Other processors changed from knpa,benj to knpa,benj,mark1

comment:38 Changed 10 years ago by knpa

  • Owner changed from knpa to mark1
  • Status changed from accepted to assigned

comment:39 Changed 10 years ago by mark1

Photography has been split up into the 7 sites. Image event file has been edited for images which failed to record a time (using time differences from sensorstat log). Images affected are:
12200323100117G3.raw
12200323100153G3.raw
12200323100154G3.raw
12200323100155G3.raw
12200323100156G3.raw
12200323100157G3.raw
12200323100194G3.raw
12200323100195G3.raw
12200323100196G3.raw
12200323100197G3.raw
12200323100198G3.raw
12200323100239G3.raw

comment:40 Changed 10 years ago by mark1

  • Description modified (diff)

Digital Photography has been dispatched to PI on 15th January 2010

comment:41 Changed 10 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Description modified (diff)

Everything delivered - Ready for archiving.

comment:42 Changed 10 years ago by mark1

  • Owner changed from mark1 to benj

comment:43 Changed 10 years ago by knpa

  • Component changed from Archiving to Processing: general

This project seems to be a bit messed up in the repository, taking archiving status off until this is sorted.

comment:44 Changed 10 years ago by knpa

  • Component changed from Processing: general to Archiving

Project now sorted out in repository (matches the workspace lay-out of one dir for each site. Raw data is in in site A).
R-synced with workspace.
Ready to begin archiving.

comment:45 Changed 10 years ago by knpa

Beginning Archiving.
Everything appears to be r-synced and delivered.

Need to include note when we send to NEODC that the layout of delivery folders varies according to how it was delivered. It seems HS and photography were delivered divided up into separate project folders for the 7 different sites (a-g) but lidar was delivered all in one folder.

comment:46 Changed 10 years ago by knpa

Doesn't appear to be any webcam data for this project

comment:47 Changed 10 years ago by knpa

Have finished preparing for archiving.
Have requested DTG archive the project (28/04/10).

comment:48 Changed 10 years ago by knpa

  • Owner changed from benj to knpa

comment:49 Changed 10 years ago by knpa

28 extra raw image files (1-28) in site A.

comment:50 Changed 10 years ago by knpa

*above is referring to RCD

comment:51 Changed 10 years ago by knpa

can't find ipas processing information for this project, i.e. gps and extract folders are empty.

comment:52 Changed 10 years ago by knpa

Have split lidar delivery into seven site structure of this project, this makes it easier to be ingested by NEODC but note lidar is no longer in the same state that it was delivered.

comment:53 Changed 10 years ago by knpa

Made available in NEODC transfer area.

comment:54 Changed 9 years ago by knpa

NEODC have confirmed that they have received, backed-up and checked this project.

comment:55 Changed 9 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Project has been moved to non-backed up space and workspace version deleted.
Closing ticket.

comment:56 Changed 3 years ago by lah

  • Cc changed from knpa,emca,mark1,sbg,peland to knpa, emca, mark1, sbg, peland

Archiving

Tidied d, e and f, and ready for upload to NEODC.

Last edited 3 years ago by lah (previous) (diff)

comment:57 Changed 3 years ago by asm

Archiving

Will start archiving of re-processed hyperspectral data.

comment:58 Changed 3 years ago by asm

Archiving

Re-processed hyperspectral data has been archived and everythin is ready to be downloaded from CEDA. Will close this ticket.

comment:59 Changed 3 years ago by asm

Archiving

Reprocessed hyperspectral has been archived and it is available from CEDA at http://browse.ceda.ac.uk/browse/neodc/arsf/2009/GB09_07/GB09_07-2009_132a_West_coast/hyperspectral-reprocessed

Plymouth 11th May 2017

Note: See TracTickets for help on using tickets.